Find the domain of the vector functions, r(t), listed below.
using interval notation.

a) r(t)=⟨ln(2t),√(t+7),1/(√(14−t))⟩
b) r(t)=⟨√(t−1),sin(1t),t^2⟩
c) r(t)=⟨ e^(−1t),t/(√(t^2−1),t^(1/3)⟩
1
Expert's answer
2020-09-22T13:53:23-0400

The domain of the vector function is the intersection of the domains of its component functions.


a) "r(t)=⟨ln(2t),\sqrt{t+7},\frac{1}{\sqrt{14−t}}⟩."

"ln(2t)" exists for "2t>0", hence "t>0" or "t \in (0, \infty);"

"\sqrt{t+7}" exists for "t+7 \geq 0", "t \geq -7" or "t \in [-7, \infty);"

"\frac{1}{\sqrt{14−t}}" exists when "14-t>0", or "t<14," "t \in (-\infty, 14)."

"(0, \infty)\bigcap [-7, \infty) \bigcap (-\infty, 14)=(0,14)."

Answer. "(0,14)."


b) "r(t)=⟨\sqrt{t−1},sin(1t),t^2⟩."

"\sqrt{t−1}" exists when "t-1 \geq0," "t \geq 1" or "t \in [1, \infty);"

"sin(1t)" exists for "t \in \R;"

"t^2" exists for "t \in \R."

"[1, \infty) \cap \R \cap \R=[1, \infty)."

Answer. "[1, \infty)."


c) "r(t)=⟨ e^{−1t},\frac{t}{\sqrt{t^2−1}},t^{1/3}⟩."

"e^{\u22121t}" exists for "t \in \R;"

"\frac{t}{\sqrt{t^2−1}}" exists when "t^2-1>0," "t^2>1," "t \in (-\infty,-1) \cup (1, \infty)."

"t^{1\/3}" exists for "t \in \R."

"((-\infty,-1) \cup (1, \infty)) \cap \R \cap \R=(-\infty,-1) \cup (1, \infty)."

Answer. "(-\infty,-1) \cup (1, \infty)."